A gas has a volume of 8 mL at a temperature of 52 °C. What is the volume of the gas at 154 °C?


1
Expert's answer
2022-01-16T20:09:02-0500

Assuming the pressure does not change, the volume of the gas at the higher temperature is


"V_f=V_i\dfrac{T_f}{T_i}=8·\dfrac{158+273}{52+273}=10.6\text{ mL}."